### v2

# You May Speak Freely: Improving the Fine-Grained Visual Recognition Capabilities of Multimodal Large Language Models with Answer Extraction

Abstract
Despite the renewed interest in zero-shot visual classification due to the rise of Multimodal Large Language Models (MLLMs), the problem of evaluating free-form responses of auto-regressive models remains a persistent challenge. Most existing works focus on language-only tasks or don’t consider Multiple Choice Questions (MCQs) beyond 5-way options, both of which are critical capabilities to solve tasks in Fine-Grained Visual Classification (FGVC) where choice counts are in the hundreds to thousands and the choices are highly related. Furthermore, in this highly multi-way MCQ setting it is not clear how to extend LLM choice extraction to retrieval-based problems, where computing probabilities over the choice set is computationally costly. In this work we investigate nlg2choice, a simple two-stage method which first asks the MLLM an open-ended question for the task with minimal constraints, then uses text-only constrained decoding to predict the most likely choice. In retrieval settings, we compute the probability of the constrained response taking that choice with an early stopping method to significantly improve throughput. Our results show improvement over a suite of seven fine-grained visual datasets when evaluating in terms of classification and retrieval, and show that this performance holds over the various ways that users of LLMs can implement tasks in natural language.
1 Introduction
Figure 1: Proposed evaluation setup of nlg2choice. Whereas previous approaches for MLLM classification use methods like probability sequences or constrained decoding, nlg2choice first prompts the MLLM for a natural language response, then performs text-only constrained decoding on the text to extract classes. Accuracy is computed as an average over multiple rewritings of prompts as a robustness measure. Correct MLLM responses are shown in green whereas incorrect responses are red.
Zero-shot visual classification has been revitalized by the introduction of Multimodal Large Language Models (MLLMs) [llava, instructblip, gpt4v, llava_next], where one can formulate classification datasets as Visual Question Answering (VQA) tasks with Multiple Choice Questions (MCQs) and a common prompt. However, fine-grained visual classification (FGVC) poses a barrier to traditional MCQ problems due to two issues: (1) a vastly higher number of choices (often hundreds to thousands) and (2) the niche concepts that each choice represents. Recent evaluations have found that state-of-the-art instruction-tuned MLLMs (e.g. LLaVA 1.5 [llava, llava_next], InstructBLIP [instructblip], GPT-4V [gpt4v]) suffer dramatic drops in accuracy when asked to name fine-grained categories. For instance, LLaVA-1.5 (13B) achieves near-perfect accuracy on coarse categories (like “bird” vs “not bird”), but only  1–2% accuracy on fine-grained species labels in iNaturalist [finer].
Figure 2: Seemingly innocuous instruction changes greatly affect MLLM performance. When testing on the gull subset of CUB200[cub_dataset], replacing one word can change the predictions of a model (left). When rewriting an instruction in a semantically equivalent way, performance varies substantially for Qwen-2.5VL (right). Each distribution comprises 15 semantically equivalent ways of writing the same instruction. These cases highlight the need to test LLMs over various equivalent inputs to ensure robustness.
Using LLMs to extract choices from pre-existing text appears to be a potential solution to this problem. Constraining LLM’s output format to ease extraction can simplify scoring but might hinder the model’s reasoning [wang2025slot, tam2024let, wang2024look], whereas allowing free-form explanations improves reasoning at the cost of harder extraction [molfese2025right, tam2024let]. Many approaches prompt a strong LLM [chiang2023vicuna, li2023alpacaeval] such as GPT4 [achiam2023gpt] to perform second-stage answer extraction. However, the works that propose LLMs as choice extractor are either focused only on language-only tasks [lyu2024beyond, wang2024my] or do not extend beyond simple lettering or numbering schemes [mmbench]. Furthermore, no significant treatment of retrieval nor the robustness of these results under input prompt variation has been considered.
In this work, we conduct a systematic analysis into constrained decoding with LLMs as answer extractors on free-form responses to fine-grained visual recognition tasks. We come to the following conclusions:
1. Answer extraction on free-form responses improves visual recognition capabilities. We find that performance improves substantially across 7 FGVC benchmarks when viewed from both classification (Tab. 2) and retrieval (Tab. 6) perspectives. It additionally improves misclassifications, with most responses being within the same genus (Tab. 3). To overcome the computational costs associated with generating probabilities over hundreds to thousands of choices, we propose an early stopping method (Sec. 3.3) which increases throughput substantially (Tab. 1).
2. Answer extraction is robust to user variation. We test robustness by generating additional semantically equivalent ways of writing an instruction (Sec. 3.1) and see that performance improvements are statistically significant when viewing user writing as a source of randomness (Fig. 3). We also show that instructions which constrict the response decrease performance (Tab. 4) and that methods which rely on forcing the LLM to generate some initial tokens to stimulate Chain-of-Thought (CoT) reasoning do not consistently increase performance (Tab. 5).
3. Constrained decoding is a reliable answer extractor without additional training. We gauge the answer extraction abilities of LLMs out-of-the-box by labeling spans and assigning answers to free-form NLG responses (Appendix A). We find that the main bottleneck is not answer extraction (Tab. 7) as studied in previous works, but a lack of extractable content from free-form responses themselves (Fig. 4). We make this data publicly available [^1] https://github.com/Racerx250/nlg2choice.
2 Related Work
Forcing MLLMs to Generate Valid Choices
The earliest works first prompted an LLM directly for a response then did regex parsing to extract choices [wang2022self], but it was shown that direct open-ended querying often underperforms on fine-grained tasks [finer]. While generating the full probability the text of each choice sequence is as an option, it remains difficult to implement in practice due to its computational cost [robinson2023leveraginglargelanguagemodels]. The most dominant choice probability generation rank the log probabilities assigned to the option IDs (e.g. ”A/B/C/D”) from the first token prediction of the model. This approach is widely adopted in many different benchmarks and LLM evaluation studies [hendrycks2020measuring, wang2024mmlu, srivastava2022beyond, liang2022holistic, santurkar2023whose].
Parallel to this, constrained decoding techniques seek to guarantee consistent outputs by guiding the generation process during sampling by constricting the output vocabulary [beurer2024guiding, liu2024we]. There are many open-source tools that implement this approach [dong2024xgrammar, GuidanceAI2023, Gerganov2023, outlines] and while these methods ensure model responses lie within the choice set, recent evaluations have shown that they decrease performance [wang2025slot, tam2024let]. Our work sits in opposition to these recent results, where we posit that constrained decoding is advantageous in the fine-grained setting where previous approaches like first token prediction fail.
MLLMs as Answer Extractors
Recent works have shown a mismatch difference between text-based outputs and token-probability metrics of LLMs on language-only tasks, revealing that a model’s first-token likelihood often fails to correspond with its generated answer [lyu2024beyond, wang2024my, wang2024look]. Evaluating a model by reading its full text answer instead of only the first-token or forced letter has been shown to improve robustness, reinforcing the value of answer extraction methods [wang2024look]. Many works use large proprietary models like GPT-4 [achiam2023gpt] to extract answers from free-form text [mmbench, chiang2023vicuna, li2023alpacaeval].
Some recent methods have addressed the problem of answer extraction directly. xFinder [yu2024xfinder] propose a lightweight model fine-tuned specifically to map an LLM’s free-form answer to the correct choice, reporting that their method outperforms a GPT-4 based extractor struggled in terms of extraction accuracy. Similarly, SLOT [wang2025slot] proposes a model-agnostic, lightweight “Structured LLM Output Transformer” that is fine-tuned to post-process any LLM’s free-form text into JSON that exactly matches an arbitrary schema, eliminating the need to modify or constrain the generating model itself. xVerify [xVerify] is an efficient answer verifier for reasoning model evaluations that extracts final answers from complex, multi-step LLM outputs and performs robust equivalence checking—trained on the VAR dataset of multi-model, multi-round–annotated question–answer pairs and achieving over 95% accuracy and F1, demonstrating strong generalizability. Contrary to these approaches the answer extraction method proposed in this work requires no additional training. We argue that answer extraction is not a critical bottleneck when using constrained decoding for fine-grained visual classification and show that LLMs perform decent answer extraction out-of-the-box on popular benchmarks.
3 Methodology
3.1 Simulating User Variation
Given the brittleness of LLMs to prompt variations (Fig. 2), we wish to test the robustness of LLMs as choice extractors by generating a range of semantically equivalent writings of prompts for a given task. We do this by starting with a small set of hand-written templates for a classification task. Letting [choice_list] be the list of classes for a given dataset separated by newlines, we define [templatei] to be equal to the following:
(Template 1) ”What is the species of bird in this image?” (Template 2) ”What is the species of bird in this image? Answer only with species name and no other text.” (Template 3) ”What is the species of bird in this image? Answer only from the following list with no other text: [choice_list]
Where the words ”species” and ”bird” are updated to suite the benchmark dataset. Given [templatei], we use o3-high [o3] to generate 14 additional prompt variations:
(Generate Variations) What are 14 semantically equivalent ways of writing the following question? Prompt: [templatei]
Next, we perform a manual check on this set of prompts for semantic equivalence to the input prompt, resulting in a total of $45$ variations for each dataset. For the remainder of the paper, accuracies reported are across these variations. For the full list of these prompts, see Appendix C.
3.2 nlg2choice for Classification
Our approach to classification involves two stages: (1) prompting the LLM for a free-form response to a question defined by Sec. 3.1 and (2) prompting the same model again to select the most likely species given the previous response. On the second step, we enforce a valid output by using constrained decoding across the choice set. Additionally, the second stage is text-only, namely it does not see the image which produced the original response. It receives the following prompt:
(Text-Only Choice) What is the most likely species of bird indicated in this response? Response: [nlg] Answer from the following: [choice_list]
Where [nlg] is the free-form response to the dataset instruction and [choice_list] is newline-separated list of class names. Similarly, ”species” and ”bird” are updated accordingly for the given dataset.
3.3 nlg2choice for Retrieval
In thresholding-based scenarios one desires a continuous score for every query. Most commonly this is done by taking the probability of "Yes" and "No" for each image, class pair $\mathbf{x}_{i}\in\mathcal{X},y_{i}\in\mathcal{Y}$, ie.:
$\hat{y}_{i}=\text{softmax}(\begin{bmatrix}p(\texttt{"Yes"}|\mathbf{x}_{i},\texttt{[prompt]})\\ p(\texttt{"No"}|\mathbf{x}_{i},\texttt{[prompt]})\end{bmatrix}$
Where $p(\cdot)$ is the next-token probability distribution of a MLLM and [prompt] describes a two-way prediction task on one class of the dataset, eg. "Is this a [cname]?" and [cname] is the name of some class. However, this requires $|\mathcal{X}|\cdot|\mathcal{Y|}$ forward passes of the model.
Method Full Prob Yes/No Truncated Prob
CUB200 687 200 47
(+0%) (+244%) (+1362%)
Flowers 257 102 12
(+0%) (+152%) (+2042%)
Aircrafts 479 100 79
(+0%) (+379%) (+506%)
Cars 1668 196 125
(+0%) (+751%) (+1234%)
Food 269 101 24
(+0%) (+166%) (+1021%)
NABirds 3294 555 701
(+0%) (+494%) (+370%)
iNat-Birds 5450 1486 480
(+0%) (+267%) (+1035%)
Table 1: Number of forward passes needed for different methods in the second-stage answer extraction of nlg2choice. ”Full Prob” refers to the probability of a full class name $p(\texttt{[cname]})$ under the LLM, where [cname] is a common name of a class. “Yes/No” refers to the common $|\mathcal{X}|\cdot|\mathcal{Y}|$ method of calculating probability over all the classes $y_{i}\in\mathcal{Y}$. “Truncated Prob” refers to the process outlined in Sec. $3.$ The percentages in the “Yes/No” and ”Truncated Prob” columns refer to the throughput increase over caculating the full sequence probabilities ”Full Prob.” Every setting for truncation is faster than “Yes/No” except for one benchmark - NABirds [nabirds_dataset]. The fastest method for each benchmark is bolded.
To improve throughput, we propose to update the second step of nlg2choice to stop generating probabilities when the remaining tokens for a given choice do not occur in any other choices. For example, take the case of "Baltimore Oriole" which can be decomposed into ["B", "altimore", " Ori", "ole"]. We can see that in the CUB200 species list there are many names which share the first token "B", ie. "Bewick Wren" or "Belted Kingfisher". However, "Baltimore Oriole" is the only choice name which has "altimore" as its second token. Thus, the probabilities for [" Ori", "ole"] are not calculated. In terms of probabilities, this is equivalent to:
$p_{full}(\texttt{"Baltimore Oriole"})=p(\texttt{"B"})$ (1)
$\hskip 56.9055pt\times p(\texttt{"altimore"}|\texttt{"B"})$
$\hskip 56.9055pt\times p(\texttt{" Ori"}|\texttt{"Baltimore"})$
$\hskip 56.9055pt\times p(\texttt{"ole"}|\texttt{"Baltimore Ori"})$
$p_{trunc}(\texttt{"Baltimore Oriole"})=p(\texttt{"B"})$ (2)
$\hskip 56.9055pt\times p(\texttt{"altimore"}|\texttt{"B"})$
$\hskip 56.9055pt\times{\color[rgb]{.5,.5,.5}\cancel{p(\texttt{" Ori"}|\texttt{"Baltimore"})}}$
$\hskip 56.9055pt\times{\color[rgb]{.5,.5,.5}\cancel{p(\texttt{"ole"}|\texttt{"Baltimore Ori"})}}$
On many datasets, this change produces substantial throughput increase as many choice names can easily be determined by their first or second tokens. The special property of truncated probability sequences is that they have the same probabilities as sampling with constrained generation. This throughput increase is depicted in Tab. 1.
4 Experiments
Datasets
We conduct experiments over a benchmark of 7 FGVC datasets: CUB [cub_dataset] (200 classes), Flowers 102 [flowers_dataset] (102 classes), Stanford Cars [cars_dataset] (196 classes), FGVC Aircrafts [aircrafts_dataset] (100 classes) and Food101 [food_dataset] (101 classes), NABirds [nabirds_dataset] (555 classes), and iNaturalist-Birds [inat_dataset] (1486 classes). We choose these benchmarks because they are well-known for fine-grained visual classification and are traditionally difficult for zero-shot visual models to solve. iNaturalist-Birds is the validation split of the iNat2021 dataset where the the ”class” field of each row is equal to ”Aves,” resulting in a test size of $14960$. For each dataset we use the full test set and the training split remains unused.
Evaluation Metrics   For the classification setting we use accuracy averaged over 15 semantically equivalent prompts, whereas for the retrieval setting we use Mean Average Precision (mAP). Choice lists are the names of classes themselves rather than letters or numbers. As compared to previous works [tan2025vision], we perform the full-way class prediction rather than subsetting the choices to a 4-way prediction. However, we provide an experiment evaluating this setting in Appendix B.
Method Birds Flowers Aircrafts Cars Foods NABirds iNat-Birds Average
Qwen-2.5VL-7B
choice 40.08 51.29 44.97 26.80 57.93 37.77 17.29 39.45
nlg2choice 47.72 63.22 59.96 37.52 68.99 44.88 21.47 49.11
(+7.64) (+11.93) (+14.99) (+10.72) (+11.06) (+7.11) (+4.18) (+9.66)
nlg2choiceopen 59.88 78.03 61.54 47.13 72.26 50.19 25.86 56.41
(+19.80) (+26.74) (+16.57) (+20.33) (+14.33) (+12.42) (+8.57) (+16.97)
Llama-3.2-Vision-11B
choice 42.27 45.68 31.38 24.68 64.70 36.69 13.82 37.03
nlg2choice 46.96 51.16 30.75 32.72 64.52 41.34 15.81 40.47
(+4.69) (+5.48) (-0.63) (+8.04) (-0.18) (+4.65) (+1.99) (+3.43)
nlg2choiceopen 49.59 62.24 37.12 36.78 72.72 46.37 13.85 45.52
(+7.32) (+16.56) (+5.74) (+12.10) (+8.02) (+9.68) (+0.03) (+8.49)
Intern3VL-8B
choice 11.27 24.68 15.84 14.78 44.12 6.90 2.28 17.12
nlg2choice 13.81 32.52 18.63 18.34 43.36 7.06 1.95 19.38
(+2.54) (+7.84) (+2.79) (+3.56) (-0.76) (+0.16) (-0.33) (+2.26)
nlg2choiceopen 18.47 37.50 21.69 22.15 51.48 12.39 4.29 24.00
(+7.20) (+12.82) (+5.85) (+7.37) (+7.36) (+5.49) (+2.01) (+6.87)
Table 2: Performance across fine-grained visual classification (FGVC) benchmarks and architectures. The “Birds” dataset refers to CUB200. The first row of each model refers to using constrained decoding directly on the model across the class list of each dataset using the prompt in the header. “choice” is performing constrained decoding on the prompt "What is the {type} of {domain} in this image? Answer with {type} only." “nlg2choice” refers to letting the model freely respond to the prompt, then using constrained decoding in a text-only approach. “nlg2choiceopen” refers to removing the appended response-style instruction, eg. "Answer with {type} only." All models are their instructed-tuned version. Best performance is bolded for each model.
Method Genus Matches % of Data
Qwen-2.5VL-7B
choice 45.99 59.92
nlg2choice 56.57 (+10.56) 52.28
nlg2choiceopen 63.90 (+17.91) 40.12
Llama-3.2-Vision-11B
choice 42.52 57.73
nlg2choice 61.71 (+19.19) 53.04
nlg2choiceopen 66.37 (+23.85) 50.94
Intern3VL-8B
choice 36.77 88.73
nlg2choice 48.79 (+12.02) 86.19
nlg2choiceopen 60.36 (+23.59) 81.53
Table 3: Accuracy of genus-level predictions on misclassified examples. ”Genus Matches” refers to the accuracy of model with respect to the translating each species to its corresponding genus. ”% of Data” refers to the amount of data misclassified by the model and is equivalent to 1accuracy1-accuracy in Tab. 2. The species to genus mapping manually created.
Implementation Details   Constrained decoding is implemented using the Outlines [^2] https://github.com/dottxt-ai/outlines library [outlines] and classes for each setting are detailed in Appendix C. We start by generating 15 prompts described in Sec. 3.1, then replacing the key features like domain (”birds,” ”flowers,” etc.) and class types (”species,” ”variant,” etc.) for each dataset. This allows us to compare directly between questions for a given model or dataset, and we also test the performance of appending class lists and common CoT methods. For our experiments, we test only open-source models. Specifically, we test the Qwen-2.5VL [qwen25vl], Llama-3.2 [llama3], and InternVL3 [internvl3] architectures. For hyperparameters, we use the recommended generation parameters as defined by the HuggingFace card and take max new tokens to be 512. Experiments were ran in a heterogeneous environment of single-GPU nodes. All nodes required a GPU with at least 23GB VRAM, 16 CPU cores, and 64GB RAM.
5 Results
5.1 Zero-Shot Classification Performance
Answer extraction is better than directly answering.
In Tab. 2 we report the performance of nlg2choice across various architectures and datasets. All architectures enjoyed a bump in performance across individual datasets and on average (+16.97, +8.49, +6.87 across Qwen-2.5VL, Llama-3.2V, and Intern3), but model-specific differences can be clearly seen. Likewise, we find that removing explicit answer formatting from the model (changing from “nlg2choice” and “nlg2choiceopen”). We find the Aves subset of iNaturalist [inat_dataset] to be the hardest among all the benchmarks, resulting in the smallest increase in performance across all models.
For model-specific differences, we find that Qwen-2.5VL performs the best on average using nlg2choice, having an average accuracy of 56.91 across the 7 datasets. We also find that Qwen-2.5VL experiences the greatest gain, having an average accuracy difference of +16.97 from the base constrained decoding method to nlg2choiceopen. Comparatively, we see that Intern3-VL performs the worst across architectures, having the overall smallest increase when using nlg2choiceopen (+6.87) and the lowest performance (24.00). However, we do still see improvements over all benchmarks.
Misclassifications are more aligned with ground truth.
A natural question is to ask how the classifications qualitatively change under nlg2choice. For an example, take Fig. 1 where the ground truth label is "Ivory Gull": under constrained decoding the model responds with "Crested Auklet" whereas a higher quality misclassificaiton would be "Herring Gull", "California Gull", …, ie. a class within the same genus as Ivory Gull. In Tab. 3 we qualitatively gauge how misclassifications change when using nlg2choice by measuring its genus-level performance. We construct a species to genus mapping on the CUB200 dataset manually, then map the species prediction in previous results to genus. Simply put, we find that nlg2choice makes more reasonable errors, switching to nlg2choice incurs a genus-level performance improvement of +17.91, +23.85, +23.59 for Qwen-2.5VL, Llama-3.2V-11B, and Intern3VL-8B.
Figure 3: Difference in accuracies on the question-level between nlg2choice and constrained decoding on CUB200. Each distribution is made up of the performance difference when switching to nlg2choice over 15 questions. nlg2choice uses an open-ended prompt whereas constrained decoding uses an instruction to increase brevity ("... Answer with species only.").
Improvements are statistically significant. Rather than ask about the average performance of different methods under user variation, we wish to quantify how much performance is gained for a specific question by switching to nlg2choice. We calculate this by taking the difference between the performance of constrained decoding and nlg2choiceopen for each question, then compute summary statistics over the set of questions. This forms a distribution of changes in accuracy, with significant results having a positive mean and 95% confidence interval. The performance difference for the various settings in outlined in Fig. 3.
We find that for the proposed settings, switching from constrained decoding with an instruction to increase brevity ("... Answer with species only.") to nlg2choiceopen comes with an average $\mu=\mathbf{18.96},\mathbf{7.38},\mathbf{7.72}$ percentage increase in accuracy for Qwen-2.5VL, Llama-3.2V-11B, and Intern3VL-8B, respectively. Additionally, we find that these increases are within the bounds of statistical significance ($\sigma=\mathbf{2.01},\mathbf{1.69},\mathbf{2.95}$).
Constrictive instructions reduce performance. A common means of steering models to produce valid choices is through appending explicit instructions onto the instruction which enumerate the available choices, typically by means of lettering or enumeration. We wish to quantify the effect of adding this explicit instruction to the model. In Tab. 4 we show the performance of constrained decoding and nlg2choice under three variations of instruction steering: (1) no explicit choice steering ("What is the {type} of {domain} in this image?"), (2) encouraging succinct prediction ("... Answer with {type} only"), and (3) explicitly enumerating choices ("... Answer from {choice_list}").
Method Birds Flowers
"What is the {type} of {domain} in this image?"
Qwen-2.5VL 0.92 2.53
+ nlg2choice 59.88 (+58.96) 78.03 (+75.5)
Llama-3.2-Vision 7.67 27.50
+ nlg2choice 49.59 (+41.92) 62.24 (+34.74)
Intern3-VL 2.28 5.01
+ nlg2choice 18.47 (+16.19) 37.5 (+32.49)
"... Answer with {type} only."
Qwen-2.5VL 40.08 51.29
+ nlg2choice 47.72 (+7.64) 63.22 (+11.93)
Llama-3.2-Vision 42.27 45.68
+ nlg2choice 46.96 (+4.69) 51.16 (+5.48)
Intern3-VL 11.27 24.68
+ nlg2choice 13.81 (+2.54) 32.52 (+7.84)
"... Answer from {choice_list}."
Qwen-2.5VL 38.71 69.19
+ nlg2choice 41.05 (+2.34) 72.48 (+3.29)
Llama-3.2-Vision 13.50 44.83
+ nlg2choice 46.68 (+33.18) 66.03 (+21.2)
Intern3-VL 22.34 35.24
+ nlg2choice 23.48 (+1.14) 35.73 (+0.49)
Table 4: Classification performance across instruction templates with varying levels of constrictiveness. The model name refers to default constrained decoding. The difference in performance is shown in parentheses. The best performing model and method for each input instruction is bolded. “nlg2choiceopen” corresponds to the top section of the table.
We see that for every template, switching from constrained decoding to nlg2choice improves performance. The most stark setting is the most open-ended prompt, where we see an average increase of +67.23, +38.33, and +24.34 for Qwen-2.5VL, Llama-3.2V, and Intern3-VL, respectively. Surprisingly, explicit choice instruction shows the weakest improvement for Qwen-2.5VL and Intern3-VL (+2.82 and +0.81) but strong improvement for Llama-3.2V (+27.19). Otherwise, keeping in line with previous results, we find that Qwen-2.5VL is the greatest benefactor of performance.
Method Birds Flowers
"What is the {type} of {domain} in this image?"
Qwen-2.5VL 59.88 78.03
Llama-3.2-Vision 49.59 62.24
Intern3-VL 18.47 37.50
... "Let’s think step by step." [kojima2022large]
Qwen-2.5VL 54.93 (-4.95) 70.16 (-7.87)
Llama-3.2-Vision 39.87 (-9.72) 39.22 (-23.02)
Intern3-VL 19.11 (+0.64) 30.82 (-6.68)
... "First," [ahn2022can]
Qwen-2.5VL 53.51 (-6.37) 72.80 (-5.23)
Llama-3.2-Vision 41.84 (-7.75) 45.31 (-16.93)
Intern3-VL 19.65 (+1.18) 33.46 (-4.04)
... "Let’s solve this problem by splitting
it into steps." [reynolds2021prompt]
Qwen-2.5VL 54.69 (-5.19) 59.56 (-18.47)
Llama-3.2-Vision 40.11 (-9.48) 38.08 (-24.16)
Intern3-VL 19.68 (+1.21) 31.06 (-6.44)
Table 5: Classification performance when using assistant output prepending CoT methods. All results are reported using nlg2choice as an answer extraction method. The proposed setting is the first section in gray. Forced assistant prepending text is shown in blue.
CoT doesn’t consistently improve performance. Next, we ask how much performance can be improved by using CoT-style instructions. We investigate methods which induce reasoning by forcing the model to generate some initial tokens that are common to thought-out reasoning (eg. "Let’s think step by step"), then letting the model freely respond. In Tab. 5 we choose three of these methods and report their performance and difference between the default nlg2choiceopen setting.
In general, we see that these methods do not consistently improve performance, resulting in an average decrease of -8.01, -15.18, and -2.36 across Qwen-2.5VL, Llama-3.2V, and Intern3-VL, respectively. The one exception to this rule is the CUB200 dataset with Intern3-VL, where we see an average increase of +1.01.
5.2 Zero-Shot Retrieval Performance
Answer extraction is better than directly answering.
In Tab. 6, we report the performance of various models when viewing FGVC datasets as a one-vs-rest retrieval task. Like the classification setting, we find that when compared to exhaustive “Yes/No” questioning, nlg2choice generally improves with the exception of the Stanford Cars [cars_dataset], in which performance across all models decreases. We again find Qwen-2.5VL to receive the greatest improvement, with +8.16 average mAP increase across the 5 datasets. Likewise, we find Intern3-VL receives the least improvement, with an average increase of +4.43.
Unlike classification, we do not find that changing the prompt to be open-ended (nlg2choiceopen, ie. removing "Answer from {species_list}" from the end of the input prompt) does not uniformly increase performance. However, we do still find that open-ended prompts generally perform better on average with nlg2choice (+8.16 vs. +6.96) and are less variable across all models.
Method Birds Flowers Aircrafts Cars Foods
Baseline
Random 0.64 1.24 1.23 0.62 1.02
Qwen-2.5VL-7B
Yes/No 51.01 61.19 51.07 77.58 82.67
nlg2choice 57.79 82.56 65.14 68.61 84.30
(+6.78) (+21.37) (+14.07) (-8.97) (+1.63)
nlg2choiceopen 59.01 86.42 64.03 69.82 85.05
(+8.00) (+25.23) (+12.96) (-7.76) (+2.38)
Llama-3.2-Vision-11B
Yes/No 41.24 63.59 34.24 66.03 78.81
nlg2choice 51.14 68.64 39.39 56.46 86.73
(+9.90) (+5.05) (+5.15) (-9.57) (+7.92)
nlg2choiceopen 52.61 74.55 40.00 59.35 88.84
(+11.37) (+10.96) (+5.76) (-6.68) (+10.03)
Intern3VL-8B
Yes/No 15.22 38.32 17.93 31.83 61.46
nlg2choice 26.97 45.13 20.47 25.81 66.55
(+11.75) (+6.81) (+2.54) (-6.02) (+5.09)
nlg2choiceopen 30.51 41.12 20.69 26.84 67.76
(+15.29) (+2.80) (+2.76) (-4.99) (+6.30)
Table 6: Performance across fine-grained visual classification (FGVC) benchmarks and architectures when interpreted as retrieval tasks. Numbers represent Mean Average Precision (mAP). Every query is a one-vs-rest task, eg. for CUB200 (Birds) the prior for a single PR curve is $1/200$.
5.3 Answer Extraction Performance
In Appendix A we detail the process of labeling data for species extraction from free-form responses. First, labelers are shown responses from various models to a dataset instruction and image, then are instructed to highlight the first incidence of the predicted species within the generated text. Then, they are instructed to confirm whether the species highlighted matches the choice selection by the model, and in the failure case to indicate the species predicted or whether there was a schema failure. This process results in a set of common outcomes, ie. “schema failure,” “no species predicted,” “answer={species}”. In Fig. 4 we show the makeup of the labeled data in terms of these various outcomes.
Models often respond out of schema. Depicted in Fig. 4, we see that 34.64% of free-form responses by models have an answer within their text that does not occur within the dataset schema. In other words, the worst-case scenario is that only 65.35% of examples are correctly classifiable. However, this turns out to not be the case due to the fact that a large percentage of these responses have real species within them (70.75% of failures) which can be resolved to the correct answer, but that the predicted species itself cannot be easily put within the given schema. We hypothesize that many of these responses are grouped into the correct answer due to either genus-level answers in the dataset, eg. CUB200’s labels include the genus "Mockingbird" which matches to "Northern Mockingbird" within the free-form text, or being a closely-related but technically different species, eg. "Glaucous-winged Gull" in schema versus "Glaucous Gull" in free-form text.
Figure 4: Breakdown of the small labeled set of natural language responses. ”Other” contains two main categories, ”Refused to answer” and ”No discernable information.” Examples of these are given in Appendix A.
Motivation of different methods. In Tab. 7 we show the performance of main architectures in terms of extraction performance over various methods when sub-setting the data to examples where an answer within the schema is clearly apparent. We include two additional variations: (1) nlg2nlg, the free-form response given the instruction to rephrase the MLLM answer and (2) nlg2nlg2choice, the constrained decoding off one extra round of rephrasing. We include nlg2nlg due to the fact that we wish to test the direct distance to the highlighted text created during labeling and include nlg2nlg2choice to gauge the distance between the interpretation of that text into an answer.
LLMs are decent answer extractors out-of-the-box. In the default nlg2choiceopen setting, we find high answer extraction success rates for Qwen-2.5VL (97.93) and Intern3-VL (93.26), but lower results for Llama-3.2V (79.02). We find that the raw text extracted by the models (”nlg2nlgexact”) seldom fits neatly into the dataset schema, having large performance drops (-16.47, 50.77, -25.91). However, the output of nlg2nlgexact retains the relevant information from the original free-form text, as indicated by the nlg2nlg2choice scores which has much smaller performance differences (+0.61, -2.72, and -4.39). We hypothesize that the nlg2nlg text still contains artifacts such as "The species is ..." or scientific names, which lower exact match performance.
Method Qwen2.5VL Llama3.2V Intern3VL Overall
xFinder [yu2024xfinder]
Qwen-1.5-0.5B [qwen1.5] 92.05 88.67 71.76 86.27
Llama-3-8B-Instruct [llama3] 80.13 80.67 75.29 79.27
Qwen-2.5VL 8B
nlg2choice 98.01 98.67 96.47 97.93
nlg2nlgexact 80.79 84.00 78.82 81.61
(-17.22) (-14.67) (-17.65) (-16.32)
nlg2nlg2choice 98.01 97.33 100.00 98.19
(+0.00) (-1.34) (+3.53) (+0.26)
Llama-3.2-Vision 11B
nlg2choice 77.48 82.00 76.47 79.02
nlg2nlgexact 31.79 26.00 25.88 28.24
(-45.69) (-56.00) (-50.59) (-50.78)
nlg2nlg2choice 76.82 80.00 70.59 76.68
(-0.66) (-2.00) (-5.88) (-2.34)
Intern3-VL 8B
nlg2choice 94.04 94.00 90.59 93.26
nlg2nlgexact 59.60 78.67 62.35 67.62
(-34.44) (-15.33) (-28.24) (-25.64)
nlg2nlg2choice 90.06 85.33 90.59 88.34
(-3.98) (-8.67) (+0.00) (-4.92)
Table 7: Performance of various MLLMs across a small labeled set. Differences are measured from the “nlg2choice“ row of each model.
6 Limitations
This work has three main limitations. First, our experiments were conducted primarily with medium-sized models (8B-11B) and we only used open-source LLMs from a few different families: Qwen, Llama, and Intern. Secondly, while nlg2choice avoids the need for specialized training data or architectural modifications and is more aligned with the true responses of the model, the method still requires significant computational resources to generate and process the textual descriptions. Lastly, while nlg2choice shows promise in classification, it still remains to be seen whether the approach scales to multi-label problems.
7 Conclusion
In this work, we presented nlg2choice, a simple approach for zero-shot fine-grained visual classification and retrieval that extracts fine-grained answers from MLLM free-form responses. Through experiments across multiple architectures and datasets, we demonstrated that nlg2choice consistently outperforms constrained decoding and exhaustive questioning methods in both classification accuracy and retrieval performance. We also validated the answer extraction directly through labeling predicted species in unstructured text, where we showed that models are capable of extraction their own answers without any labeled data.
Appendix A Labeling Species Answer Extraction Data
Motivation.
For labeling answer extraction data, we prepare a user interface built with Label Studio which shown in Fig. 5. On the right contains information on the inputs to the model, specifically the question, image, and ground truth label. On the left is main interface, which consists a free-form response from a model to the example displayed on the right, as well as the choice predicted by the model via nlg2choice. Labeling consists of two pieces: (1) identification and highlighting of text spans containing answers to the input question and (2) matching the highlighted text span to a class within the given schema.
Figure 5: Labeling interface for efficacy of the answer extraction stage of nlg2choice. Labelers are asked to highlight the name predicted in free-form responses from various models and whether the name occurs easily within the dataset choice list. Above shows an interesting example: for the Chipping Sparrow on the right, the model freely responded with ”Painted Bunting” and answer extraction made an additional mistake by predicting ”Scarlet Tanager.”
Text span labeling.
For the text span labeling task, labelers read through the text to identify the first incidence of a predicted species. Before highlighting the species found within the text, the prediction is confirmed to be a real species through a brief internet search. If there are any issues with highlighting the predicted species, a text description of the issue is written in the input text box at the bottom and the example is skipped.
Choice assignment.
Once the span is labeled, we next wish to either assign the span to a class within the dataset or tag the span as a schema failure. For expediency, the labelers are also equipped with the nlg2choice prediction and a tool for sorting most like class names with fuzzy matching score. When the nlg2choice predict matches the text span, the prediction is ”Consistent with nlg” and the labelers check the button. Otherwise, ”Inconsistent with nlg” is checked and a natural language explanation is put in the bottom text box. When the nlg2choice is inconsistent, “answer={species}” is put within that text box.
Appendix B Zero-Shot 4-Way VQA Performance
Method
CUB200
iNat-Animal
iNat-Plant
ImgNet-Animal
ImgNet-Artifact
Qwen-2.5VL-7B
A/B/C/D 65.50 41.33 41.61 85.20 80.01
nlg2choiceopen 67.43 40.19 40.03 84.03 81.81
(+1.93) (-1.14) (-1.58) (-1.17) (+1.80)
Llama-3.2-Vision-11B
A/B/C/D 65.52 32.44 31.88 79.93 75.89
nlg2choiceopen 68.79 31.52 31.57 81.39 75.29
(+3.27) (-0.92) (-0.31) (+1.46) (-0.60)
Intern3VL-8B
A/B/C/D 50.52 35.40 36.39 77.50 69.41
nlg2choiceopen 47.30 35.86 35.64 81.91 70.02
(-3.22) (+0.46) (-0.75) (+4.41) (+0.61)
Table 8: 4-way VQA accuracy across fine-grained classifcation tasks.
Answer extraction does not consistently increase performance.
Following recent work [tan2025vision], we evalute nlg2choice on popular fine-graiend classification datasets as a 4-way VQA task. iNat-Animal and iNat-Plant are the rows of iNat21 [inat_dataset] when subset to rows which whose kingdom matches ”Animalia” and ”Plantae.” ImgNet-Animal and ImgNet-Artifact are the rows of ImageNet-1K [imagenet_dataset] subset to rows which have an anceter of ”Animal” and ”Artifact” in the WordNet hierarchy [wordnet_dataset]. The choices for each example are from the top 3 SigLIP [zhai2023sigmoid] predictions and the correct class. For the nlg2choice prompt we use the same prompt as the previous work. We report the open-ended question performance in Tab. 8 where we find that it slightly underperforms or has not effect on the lettering approach. Specifically, we find that accuracy changes by -0.03, +0.58, and +0.30 on average for Qwen-2.5VL, Llama-3.2V, and Intern3VL, respectively.
Decreasing the choice set improves performance. We are also able to compare the 4-way VQA setting of CUB200 directly to the many-way prediction displayed in Tab. 2. We see that subsetting the full choice set to four choices incurs a performance improvement of +7.55, +19.20, and +28.83 for Qwen-2.5VL, Llama-3.2V, and Intern3VL, respectively.
Appendix C Prompt Variations
"What { type } is this { domain }?" "What is the { type } of this { domain }?" "What is the { type } of the { domain }?" "What is the { type } of the { domain } in this image?" "What is the { type } of the { domain } in the image?" "Identify this { domain }’s { type }." "Name the { type } shown in the image." "Which { domain } { type } is pictured here?" "Classify the { type } of this { domain }." "What { domain } { type } does the photo depict?" "Determine the { type } of the { domain } in view." "Provide the common name of this { domain }." "To which { type } does this { domain } belong?" "Label the { type } of the { domain } shown." "Recognize and state this { domain }’s { type }."
Table 9: Base template variations generated by Sec. 3.1
Dataset Type Domain
CUB200 species bird
Flowers species flower
Aircrafts variant aircraft
Cars year, make, and model car
Foods name food
NABirds species bird
iNat-Birds species bird
Table 10: Dataset variables for filling out prompt templates.